# TTPXHunter: Actionable Threat Intelligence Extraction as TTPs form Finished Cyber Threat Reports

### 05 Mar 2024 v1

@@ Abstract
Understanding the modus operandi of adversaries aids organizations in employing efficient defensive strategies and sharing intelligence in the community. This knowledge is often present in unstructured natural language text within threat analysis reports. A translation tool is needed to interpret the modus operandi explained in the sentences of the threat report and translate it into a structured format. This research introduces a methodology named TTPXHunter for the automated extraction of threat intelligence in terms of Tactics, Techniques, and Procedures (TTPs) from finished cyber threat reports. It leverages cyber domain-specific state-of-the-art natural language processing (NLP) to augment sentences for minority class TTPs and refine pinpointing the TTPs in threat analysis reports significantly. The knowledge of threat intelligence in terms of TTPs is essential for comprehensively understanding cyber threats and enhancing detection and mitigation strategies. We create two datasets: an augmented sentence-TTP dataset of $39,296$ samples and a $149$ real-world cyber threat intelligence report-to-TTP dataset. Further, we evaluate TTPXHunter on the augmented sentence dataset and the cyber threat reports. The TTPXHunter achieves the highest performance of $92.42\%$ f1-score on the augmented dataset, and it also outperforms existing state-of-the-art solutions in TTP extraction by achieving an f1-score of $97.09\%$ when evaluated over the report dataset. TTPXHunter significantly improves cybersecurity threat intelligence by offering quick, actionable insights into attacker behaviors. This advancement automates threat intelligence analysis, providing a crucial tool for cybersecurity professionals fighting cyber threats.
@@
Threat Intelligence, TTP Extraction, MITRE ATT&CK, Natural Language Processing, Domain-specific Language model, TTP Classification
@@ I Introduction
In the ever-evolving landscape of cybersecurity, Advanced Persistent Threats (APTs) represent a significant challenge to worldwide security. Countering APTs requires the development of sophisticated measures, which depend on the detailed extraction and analysis of threat intelligence related to APTs [20]. It involves delving into the attacker’s modus operandi in terms of Tactics, Techniques, and Procedures (TTPs)  [^1] Categorized in the MITRE ATT&CK Framework [6] explained in the threat reports released by security firms [19]. These reports, written in unstructured natural language, describe cyber adversaries’ modus operandi. Converting this information into a machine-readable structured format improves threat intelligence efforts and is crucial for comprehending and mitigating potential threats.
Extracting TTPs from such reports is also crucial for recommending defensive mechanisms. However, this process often encounters challenges, including a lack of publicly available structured data, difficulties posed by polymorphic words, and the challenge of interpreting the contextual meanings of sentences present in the threat report and mapping them to MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) TTP [21]. We combat these challenges in one of our previous works, which presents a TTP extraction tool named TTPHunter [1]. It can automatically identify and catalog $50$ most frequently used distinct TTPs with the f1-score of $0.88$. TTPHunter focused on top 50 TTPs present in the ATT&CK framework because of the limited data availability for the remainder of the TTP class. The remainder of the TTPs are particularly those that are either newly emerging or less commonly used. This gap undermines the effectiveness of TTPHunter, as incomplete TTP intelligence can lead to a reactive rather than proactive security posture. Addressing this gap by advancing TTPHunter is crucial to ensure that threat intelligence is exhaustive and reflects the latest adversarial tactics.
@@
No. of words sentences present in the report
@@
List of augmented sentence for the sentence $S$
@@
Top-5 selected word to create augmented sentence
@@
Therefore, This study presents TTPXHunter, an extended version of TTPHunter, which meticulously refines and expands to recognize an impressive array of $193$ TTPs. We address the limited data problem of TTPHunter by introducing an advanced data augmentation method. This method meticulously preserves contextual integrity and enriches our training dataset with additional examples for emerging and less commonly used TTPs. Moreover, we also employ a domain-specific language model that is finely tuned to grasp the nuanced, context-driven meanings within this domain and enhance both the augmentation process and the TTP classification. This advancement significantly broadens the scope of threat intelligence gleaned from threat reports and offers deeper insights into the TTPs employed by cyber adversaries in APT campaigns. By extending TTPHunter’s extraction capabilities, this research contributes to the critical task of threat intelligence gathering, providing security analysts and practitioners with a more comprehensive toolset for identifying, understanding, and countering APTs.
TTPXHunter applies to a threat report denoted as $\mathcal{R}$ and extracts set of TTPs denoted as $\hat{\mathcal{T}}$ explained within the report, where $\hat{\mathcal{T}}\subseteq\mathcal{T}$ and $\mathcal{T}=\{t_{1},t_{2},t_{3},\ldots,t_{m}\}$ which denotes a set of target labels having $m$ number of TTPs. The methodology initiates with the transformation of sentences segmented from threat report, denoted as =$\mathcal{R}=\{S_{1},S_{2},\ldots,S_{n}\}$; $n$ represents number of sentences in the report, into a high-dimensional feature space, i.e., $768-dimension$. This transformation, achieved through a cyber domain-specific embedding function $f(.)$, maps each segmented sentence $S_{i}$ to a $768-dimension$ feature vector $v_{i}$ as
@@
The embedding function encapsulates cybersecurity language’s rich semantic and syntactic properties, setting the stage for nuanced inference. TTPXHunter exhibits an inferential capability to predict a specific TTP for each segmented sentence based on its embedded feature vector:
@@
  • We introduce TTPXHunter, an extended version of TTPHunter [1], which extracts TTPs from threat intelligence reports using the SecureBERT [2] language model. We fine-tune the model on our prepared augmented sentence-TTP dataset.
  • We introduce a data augmentation method that utilizes the cyber-domain-specific language model. This approach creates sentences related to TTPs by preserving their contextual meaning while preparing new, varied sentences.
@@
  • We perform an extensive evaluation to measure the efficiency of TTPXHunter over two different types of datasets: The augmented sentence dataset and the Threat report dataset. For the report dataset, we prepare the ground truth manually.
@@ II Related Work
Extract attack phases using sequence tagging model and mapping these patterns to TTP using cosine similarity
@@
Initially, Husari et al. [4] present TTPDrill, which extracts threat actions based on ontology and matches TTP’s knowledge base with extracted threat action using the BM25 matching technique. Legoy et al. [9] present rcATT, a classification tool based on Machine Learning (ML) algorithms. They use Term Frequency–Inverse Document Frequency (TF-IDF) to prepare the dataset and perform multi-class classification for target labels as TTP. The recast is based only on sentence keywords, as TF-IDF ignores the sentence’s word sequence and context. The model may fail when synonym words are present and polymorphic nature words (the same words have a different meaning in a different context). Li. et al. [5] introduce AttacKG, a graph template matching technique. The method obtains IOCs and constructs entity-based dependency graphs for every TTP present in ATT&CK by leveraging descriptions on the MITRE website. Further, they extract TTP from the threat report by creating an entity-based dependency graph for every report and matching it with the TTP’s templates prepared from MITRE website data. To match the templates, they use the graph alignment algorithm. AttacKG struggles to capture techniques identified by adjectives (properties of IOCs) present in the sentences rather than verbs (threat actions), such as masqueraded identity and obfuscated malware. Alam et al.  [3] introduce LADDER, a framework designed to enhance cyber threat intelligence (CTI) by extracting and analyzing attack patterns from CTI reports, addressing the limitations of traditional CTI that focuses on static indicators like IP addresses. LADDER contains a subpart named TTPClassifier, which is structured into three key steps: identifying sentences with attack pattern descriptions using a binary classification model, pinpointing and extracting these attack phrases with a sequence tagging model, and finally, mapping these patterns to TTP IDs via cosine similarity method that uses computed embedding. To deal with the polymorphic nature of words and leverage contextual information in the report sentences, Rani et al. [1] present TTPHunter. This tool leverages the language models BERT and RoBERTa to understand the context of the sentences present in the dataset and map it to the correct TTP ID present in the dataset. Due to a dearth of sentence datasets for many TTPs, TTPHunter is trained for only $50$ sets of TTPs and has yet to map the full spectrum of TTPs present in the MITRE knowledge base. After that, MITRE also introduced a TTP extraction tool named TRAM (Threat Report ATT&CK Mapping) [7]. It is based on the scientific BERT model, named SciBERT [18], a fine-tuned BERT model on a collection of scientific reports.
@@
In this paper, we extend our recent work TTPHunter [1] to extract TTPs present in the collected threat report. The TTPHunter is fine-tuned on traditional BERT and RoBERTa models, whereas we fine-tune cyber-domain-specific language models to identify TTPs. Aghaei et al. [2] show that the domain-specific language model can perform better than the traditional model trained on general English sentences. Domain-specific words such as Windows and registry have different meanings in terms of general and cybersecurity usage. In addition, we solve the limited dataset problem pointed out by [1] using the data augmentation method. Our proposed method can extract the full spectrum of TTPs in the MITRE ATT&CK knowledgebase. A comparison of our proposed model TTPXHunter with the literature is shown in Table II.
@@ III-A MITRE ATT&CK Framework
The MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) Framework [6] is a globally acknowledged comprehensive knowledge base designed to understand cyber threats and adversary behaviors across different stages of a cyber attack. It introduces a standardized lexicon and framework for the classification and detailed description of attackers’ tactics, techniques, and procedures (TTPs). The ATT&CK framework comprises three primary components: Tactics, Techniques, and Procedures. Tactics represent the objectives or goals adversaries aim to achieve throughout an attack, encompassing $14$ distinct categories such as Reconnaissance, Resource Development, Initial Access, Execution, and others [6]. Techniques detail the specific methods adversaries employ to fulfill these tactics, each with a unique identifier and a comprehensive description. Examples include phishing, password spraying, remote code execution, and PowerShell exploitation. Procedures delve into the detailed execution of techniques by attackers to achieve their tactical objectives, illustrating how specific methods are applied in practice, such as deploying a spearphishing email with malicious attachments to achieve initial access. Beyond cataloging techniques, the ATT&CK framework maps real-world attack scenarios to known techniques that enable organizations to understand better how adversaries navigate through the stages of an attack. This knowledge fosters the development of proactive defenses, enhances incident response efforts, and enriches threat intelligence.
@@ III-C TTPHunter
A recent tool, TTPHunter, developed by Rani et al. [1], leverages the power of the language model BERT to extract threat intelligence in the form of TTPs from natural language threat report texts. The authors of TTPHunter fine-tune a pre-trained BERT model using a sentence-TTP dataset collected from the MITRE knowledge base to let the model understand the context of TTPs present in the MITRE ATT&CK matrix and map a given sentence to relevant TTPs. One noteworthy feature of TTPHunter is its ability to discern the significance of sentences. Rather than indiscriminately mapping all sentences from a threat report, TTPHunter selectively identifies and considers only those sentences that truly explain TTPs, disregarding irrelevant information. We implement a filtering mechanism on the model’s predicted probabilities to achieve this. Our experiments discovered that sentences with a probability higher than $0.64$ are relevant, while those below the threshold were considered irrelevant. We adopt the same threshold for identifying relevant sentences by TTPXHunter.
@@ IV Proposed Methodology: TTPXHunter
In this study, we introduce TTPXHunter, an extended version of TTPHunter, designed to overcome the challenges of limited data for less commonly encountered TTPs and improve the performance of TTPHunter. TTPXHunter incorporates a novel methodology that encompasses both data augmentation and domain-specific language models to enhance the performance of TTP classification. The notations used to explain the methodology are listed in Table. I.
TTPXHunter follows the idea used in TTPHunter of leveraging the natural language model to extract TTPs from the threat reports. TTPHunter is based on the model fine-tuned on general English sentences. The model trained on general sentences can mislead the contextual embedding for domain-specific words having different contexts comparatively. Cyber domain-specific words such as Windows and Registry have entirely different meanings in the cyber domain. Hence, we observe the need for a TTP extraction tool based on domain-specific language models to provide more accurate contextual embedding. To fill the gap of domain-specific knowledge, we leverage the SecureBERT [2] to fine-tune the proposed TTPXHunter and map sentences to the relevant TTPs they represent. TTPXHunter incorporates a filtration mechanism to identify and exclude irrelevant sentences from the TTP extraction results obtained from the threat reports. The filtering mechanism ensures the model extracts the most pertinent and meaningful TTPs for further analysis and discards unrelated sentences. The overview and architecture of TTPXHunter are shown in Fig. 1.
The TTPHunter [1] is limited to 50 prominent TTPs only out of $193$ TTPs due to the limited sentence in the database for the remainder of the TTP class. The limitations in extracting the full spectrum of TTPs lead to incomplete threat intelligence, undermining cybersecurity defenses by reducing operational efficiency and increasing the risk of security oversights. Enhancing its extraction capabilities is crucial for developing comprehensive and proactive security strategies. Therefore, we address this problem by presenting a data augmentation method in TTPXHunter. This method creates more samples for the minority TTP class.

### 21 Mar 2024 v3

# TTPXHunter: Actionable Threat Intelligence Extraction as TTPs from Finished Cyber Threat Reports

@@ Abstract
Understanding the modus operandi of adversaries aids organizations to employ efficient defensive strategies and share intelligence in the community. This knowledge is often present in unstructured natural language text within threat analysis reports. A translation tool is needed to interpret the modus operandi explained in the sentences of the threat report and convert it into a structured format. This research introduces a methodology named TTPXHunter for automated extraction of threat intelligence in terms of Tactics, Techniques, and Procedures (TTPs) from finished cyber threat reports. It leverages cyber domain-specific state-of-the-art natural language model to augment sentences for minority class TTPs and refine pinpointing the TTPs in threat analysis reports significantly. We create two datasets: an augmented sentence-TTP dataset of $39,296$ sentence samples and a $149$ real-world cyber threat intelligence report-to-TTP dataset. Further, we evaluate TTPXHunter on the augmented sentence and report datasets. The TTPXHunter achieves the highest performance of $92.42\%$ f1-score on the augmented dataset, and it also outperforms existing state-of-the-art TTP extraction method by achieving an f1-score of $97.09\%$ when evaluated over the report dataset. TTPXHunter significantly improves cybersecurity threat intelligence by offering quick, actionable insights into attacker behaviors. This advancement automates threat intelligence analysis and provides a crucial tool for cybersecurity professionals to combat cyber threats.
@@
Threat Intelligence, TTP Extraction, MITRE ATT&CK, Natural Language Processing, Domain-specific Language model, TTP Classification, Cybersecurity
@@ I Introduction
In the ever-evolving landscape of cybersecurity, Advanced Persistent Threats (APTs) represent a significant challenge to worldwide security. Countering APTs requires the development of sophisticated measures, which depend on the detailed extraction and analysis of threat intelligence related to APTs [24, 27, 31]. It involves delving into the attacker’s modus operandi in terms of Tactics, Techniques, and Procedures (TTPs)  [^1] Categorized in the MITRE ATT&CK Framework [7] explained in the threat reports, blogs, bulletins released by security firms [1, 23]. These reports, written in unstructured natural language, describe cyber adversaries’ modus operandi. Converting this information into a machine-readable structured format improves threat intelligence efforts and is crucial for comprehending and mitigating potential threats [30].
Extracting TTPs from such reports is also crucial for recommending defensive mechanisms. However, this process often encounters challenges, including a lack of publicly available structured data, difficulties posed by polymorphic words, and the challenge of interpreting the contextual meanings of sentences present in the threat report and mapping them to MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) TTP [2, 25]. We combat these challenges in one of our previous works, which presents a TTP extraction tool named TTPHunter [2]. It can automatically identify and catalog TTPs with the f1-score of $0.88$ [2]. TTPHunter focused on the top 50 TTPs in the ATT&CK framework because of the limited data availability for the remainder of the TTP class. The remainder of the TTPs are particularly those that are either newly emerging or less commonly used. This gap undermines the effectiveness of TTPHunter, as incomplete TTP intelligence can lead to a reactive rather than proactive security posture. Addressing this gap by advancing TTPHunter is crucial to ensure that threat intelligence is exhaustive and reflects the diverse adversarial tactics.
@@
No. of sentences present in the report
@@
List of augmented sentences for the sentence $S$
@@
Top-5 selected words to create augmented sentence
@@
Therefore, this study presents TTPXHunter, an extended version of TTPHunter [2], which meticulously refines and expands to recognize an impressive array of $193$ TTPs. We address the limited data problem of TTPHunter by introducing an advanced data augmentation method. This method meticulously preserves contextual integrity and enriches our training dataset with additional examples for emerging and less commonly used TTPs. Moreover, we also employ a domain-specific language model that is finely tuned to grasp the nuanced, context-driven meanings within this domain and enhance both the augmentation process and the TTP classification. Further, we also enable TTPXHunter to convert the extracted TTPs into the machine-readable format, i.e., STIX, which facilitates the automated exchange, easier analysis, and integration across different security tools and platforms [18, 34]. This advancement significantly broadens the scope of threat intelligence gleaned from threat reports and offers deeper insights into the TTPs employed by cyber adversaries in APT campaigns. By extending TTPHunter’s extraction capabilities, this research contributes to the critical task of threat intelligence gathering, providing security analysts and practitioners with a more comprehensive toolset for identifying, understanding, and countering APTs. The notations and their descriptions used to explain the methodology are listed in Table. I.
TTPXHunter applies to a threat report denoted as $\mathcal{R}$ and extracts set of TTPs denoted as $\hat{\mathcal{T}}$ explained within the report, where $\hat{\mathcal{T}}\subseteq\mathcal{T}$ and $\mathcal{T}=\{t_{1},t_{2},t_{3},\ldots,t_{m}\}$ which denotes a set of target labels having $m$ number of TTPs. The methodology initiates with the transformation of sentences segmented from threat report, denoted as =$\mathcal{R}=\{S_{1},S_{2},\ldots,S_{n}\}$; $n$ represents the number of sentences in the report, into a high-dimensional feature space, i.e., $768$-dimension. This transformation, achieved through a cyber domain-specific embedding function $f(.)$, maps each segmented sentence $S_{i}$ to a $768$-dimension feature vector $v_{i}$ as
@@
The embedding function encapsulates cybersecurity language’s rich semantic and syntactic properties and the stage for nuanced inference. TTPXHunter exhibits an inferential capability to predict a specific TTP for each segmented sentence based on its embedded feature vector:
@@
  • We introduce TTPXHunter [^2] We plan to include the source code link in the camera-ready version., an extended version of TTPHunter [2], which extracts TTPs from threat intelligence reports using the SecureBERT [3] language model. We fine-tune the model on our prepared augmented sentence-TTP dataset and convert the extracted TTPs into structured and machine-readable STIX format.
  • We introduce a data augmentation method that utilizes the cyber-domain-specific language model. This approach creates sentences related to TTPs by preserving their contextual meaning while preparing new and varied sentences.
@@
  • We perform an extensive evaluation to measure the efficiency of TTPXHunter over two different types of datasets: The augmented sentence dataset and the Threat report dataset. We manually label $149$ real-word threat reports for the report dataset.
@@ II Related Work
Extract attack phases using a sequence tagging model and map these patterns to TTP using cosine similarity
@@
Initially, Husari et al. [5] present TTPDrill, which extracts threat actions based on ontology and matches TTP’s knowledge base with extracted threat action using the BM25 matching technique. Legoy et al. [10] present rcATT, a classification tool based on Machine Learning (ML) algorithms. They use Term Frequency–Inverse Document Frequency (TF-IDF) to prepare the dataset and perform multi-class classification for target labels as TTP. The rcATT is based only on sentence keywords, as TF-IDF ignores the sentence’s word sequence and context. The model may fail when synonym words are present and polymorphic nature words (the same words have a different meaning in a different context). Li. et al. [6] introduce AttacKG, a graph template matching technique. The method obtains IOCs and constructs entity-based dependency graphs for every TTP present in ATT&CK by leveraging descriptions on the MITRE website and matching them with the TTP’s templates prepared from MITRE website data. To match the templates, they use the graph alignment algorithm. AttacKG struggles to capture techniques identified by adjectives (properties of IOCs) present in the sentences rather than verbs (threat actions), such as masqueraded identity and obfuscated malware. Alam et al.  [4] introduce LADDER, a framework designed to enhance cyber threat intelligence (CTI) by extracting and analyzing attack patterns from CTI reports, addressing the limitations of traditional CTI that focuses on static indicators like IP addresses. LADDER contains a subpart named TTPClassifier, which is structured into three key steps: identifying sentences with attack pattern descriptions using a binary classification model, pinpointing and extracting these attack phrases with a sequence tagging model, and finally, mapping these patterns to TTP IDs via cosine similarity method. To deal with the polymorphic nature of words and leverage contextual information in the report sentences, Rani et al. [2] present TTPHunter. This tool leverages the language models BERT and RoBERTa to understand the context of the sentences present in the dataset and map it to the correct TTP ID present in the dataset. Due to a dearth of sentence datasets for many TTPs, TTPHunter is trained for only $50$ sets of TTPs and has yet to map the full spectrum of TTPs present in the MITRE knowledge base. After that, MITRE also introduced a TTP extraction tool named TRAM (Threat Report ATT&CK Mapping) [8]. It is based on the scientific BERT model, named SciBERT [22], a fine-tuned BERT model on a collection of scientific reports.
@@
In this paper, we extend our recent work TTPHunter [2] to extract TTPs present in the collected threat report. The TTPHunter is fine-tuned on traditional BERT and RoBERTa models, whereas we fine-tune cyber-domain-specific language models to identify TTPs. Aghaei et al. [3] show that the cyber-domain-specific language model can perform better than the traditional model trained on general English sentences. Domain-specific words such as Windows and registry have different meanings regarding general and cybersecurity usage. In addition, we solve the limited dataset problem pointed out by [2] using the data augmentation method. Our proposed method can extract the full spectrum of TTPs in the MITRE ATT&CK knowledgebase. A comparison of our proposed model TTPXHunter with the literature is shown in Table II.
@@ III-A MITRE ATT&CK Framework
The MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) Framework [7] is a globally acknowledged comprehensive knowledge base designed to understand cyber threats and adversary behaviors across different stages of a cyber attack. It introduces a standardized lexicon and framework for the classification and detailed description of attackers’ tactics, techniques, and procedures (TTPs). The ATT&CK framework comprises three primary components: Tactics, Techniques, and Procedures. Tactics represent the objectives or goals adversaries aim to achieve throughout an attack, encompassing $14$ distinct categories such as Reconnaissance, Resource Development, Initial Access, Execution, and others [7]. Techniques detail the specific methods adversaries employ to fulfill these tactics, each with a unique identifier and a comprehensive description. Examples include phishing, password spraying, remote code execution, and PowerShell exploitation. Procedures delve into the detailed execution of techniques by attackers to achieve their tactical objectives, illustrating how specific methods are applied in practice, such as deploying a spearphishing email with malicious attachments to achieve initial access. Beyond cataloging techniques, the ATT&CK framework maps real-world attack scenarios to known techniques that enable organizations to understand better how adversaries navigate through the stages of an attack. This knowledge fosters the development of proactive defenses and offensive security practices, enhances incident response efforts, and enriches threat intelligence [23, 32].
@@ III-C TTPHunter
Our recent tool, TTPHunter [2], leverages the power of the language model BERT to extract threat intelligence in the form of TTPs from natural language threat report texts. The authors of TTPHunter fine-tune a pre-trained BERT model using a sentence-TTP dataset collected from the MITRE knowledge base to let the model understand the context of TTPs present in the MITRE ATT&CK matrix and map a given sentence to relevant TTPs. One noteworthy feature of TTPHunter is its ability to discern the significance of sentences. Rather than indiscriminately mapping all sentences from a threat report, TTPHunter selectively identifies and considers only those sentences that truly explain TTPs, disregarding irrelevant information. To achieve this, we implement a filtering mechanism on the model’s predicted probabilities. Our experiments discovered that sentences with a probability higher than $0.64$ are relevant, while those below the threshold were considered irrelevant. We adopt the same threshold for identifying relevant sentences by TTPXHunter.
@@ IV Proposed Methodology: TTPXHunter
In this study, we introduce TTPXHunter, an extended version of TTPHunter, designed to overcome the challenges of limited data for less commonly encountered TTPs and improve the performance of TTPHunter. It incorporates a novel methodology encompassing data augmentation and domain-specific language models to enhance the performance of TTP classification. The notations used to explain the methodology are listed in Table. I.
TTPXHunter follows the idea used in TTPHunter of leveraging the natural language model to extract TTPs from threat reports. TTPHunter is based on the model, which is fine-tuned on general English sentences. The model trained on general sentences can mislead the contextual embedding for domain-specific words having different contexts comparatively. Cyber domain-specific words such as Windows and Registry have entirely different meanings in the cyber domain. Hence, we observe the need for a TTP extraction tool based on domain-specific language models to provide more accurate contextual embeddings. To fill this gap of domain-specific knowledge, we leverage the SecureBERT [3] to fine-tune the proposed TTPXHunter and map sentences to the relevant TTPs they represent. TTPXHunter incorporates a filtration mechanism to identify and exclude irrelevant sentences from the TTP extraction results obtained from the threat reports. The filtering mechanism ensures the model extracts the most pertinent and meaningful TTPs for further analysis and discards unrelated sentences. The overview and architecture of TTPXHunter are shown in Fig. 1.
The TTPHunter [2] is limited to $50$ prominent TTPs only out of $193$ TTPs due to the limited sentence in the database for the remainder of the TTP class, which results in incomplete or limited threat intelligence. Enhancing its extraction capabilities is crucial for developing comprehensive and proactive security strategies by leveraging the full spectrum of threat intelligence. Therefore, we address this problem by presenting a data augmentation method in TTPXHunter. This method creates more samples for the minority TTP class.